Solve the following system by graphing
2x-y=4
2x-y=6

Answer by jainenderkapoor(61)


2x-y=4 --- (1)
2x-y=6 --- (2)
For (1) equation ---
we get
For x = 0, 0 - y = 4 or y = -4 so Point is (0, -4)
For y = 0, 2x - 0= 4 or 2x = 4 or x = 2 so point is (2, 0)
For (2) equation ---
we get
For x = 0, 0 - y = 6 or y = -6 so Point is (0, -6)
For y = 0, 2x - 0= 6 or 2x = 6 or x = 3 so point is (3, 0)
Now plot the points for both equations and draw straight lines joining the points...
You will get two lines which are parallel lines.. and never intersect..
Which shows that the system of linear equations is inconsistent (no solution exists).
Also if I look at the equations and see the coefficients of (x), of (y) and the independent term of both equations..
I get 2/2 = -1/-1 which is not equal to the independent terms 4/6
so condition for inconsistency is a1/a2 = b1/b2 not equal to c1/c2
Hope the solution is clear to you. In case you have any doubt, you are welcome to contact me at jainederkapoor@gmail.com I am also available for Online Tutoring at very reasonable price.
Jainender Kapoor